# and extensions of the log-Sobolev and Talagrand inequalities

### 04 Mar 2024 v1

@@ 2.2 The Talagrand inequality and flow’s exponential convergence
In this case, Theorem 2 reduces to the well-known result stating that the QG condition holds whenever the PŁI (21) holds; see [21, Theorem 2].
@@ 2.3 Strongly log-concave models
Any model satisfying Assumptions 1(iii) and 4 satisfies the xLSI.
@@
Given that (11) with $q=\pi_{\theta}$ reduces to the PŁI (21) with $f(\theta)=-\log(Z_{\theta})$, Theorem 4 extends the well-known fact that the PŁI holds for strongly convex $f$; e.g., see [21, Theorem 2].
@@
Under Assumption 4, the marginal likelihood has a unique maximizer $\theta_{*}$; e.g., see [23, Theorem 4 in Appendix B.3]. Hence, Corollary 3 and (24) tell us that the flow converges exponentially fast to $(\theta_{*},\pi_{\theta_{*}})$:
@@
If Assumptions 1(iii), 24 hold and $(\theta_{0},q_{0})$ belongs to $\mathcal{M}_{2}$, then
@@ 3 Error bounds for particle gradient descent
If Assumptions 45 hold and $(\theta,X)$ belongs to $\mathbb{R}^{d_{\theta}}\times\mathcal{L}^{2}(\mathbb{R}^{d_{x}})$, then the SDE (6) has an unique strong solution $(\theta_{t},X_{t})_{t\geq 0}$ such that $(\theta_{0},X_{0})=(\theta,X)$. Moreover, $(\theta_{t},\textup{Law}(X_{t}))_{t\geq 0}$ is a weak solution of the gradient flow (5) in the sense of Assumption 2.
@@
Under Assumptions 1(iii) and 35, if $X_{0}^{1,h},\dots,X_{0}^{N,h}$ are drawn independently from a distribution $q_{0}$ in $\mathcal{P}_{2}(\mathbb{R}^{d_{x}})$ and $h\leq 1/(\lambda+L)$, then
$\mathsf{d}((\Theta_{K}^{N,h},Q_{K}^{N,h}),(\theta_{*},Q_{*}^{N}))\leq\sqrt{h}A_{0,h}+\frac{L\sqrt{2}}{\lambda\sqrt{N}}\sqrt{B_{0}+\frac{d_{x}}{\lambda}}+\mathsf{d}((\theta_{0},q_{0}),(\theta_{*},\pi_{\theta_{*}}))e^{-h\lambda K}$ (27)
@@
We postpone the proof until Section 3.1 below. In short, the Theorem 7 gives us the bound (9) and an explicit expression for the proportionality constant as a function of the algorithm’s parameters, the model’s Lipschitz and concavity constants, the initial condition, and the dimension of the latent space. It follows that in order to achieve an error of $\mathcal{O}(\epsilon)$, we need only choose the step size $h$ proportional to $\epsilon^{2}$, the particle number $N$ proportional to $\epsilon^{-2}$, and the step number $K$ proportional to $\epsilon^{-2}\log(\epsilon^{-1})$. The corresponding computational cost is $\mathcal{O}(\epsilon^{-4}\log(\epsilon^{-1})[\text{eval. cost of }\ell])$; and we can lower the running time to $\mathcal{O}(\epsilon^{-2}\log(\epsilon^{-1})[\text{eval. cost of }\ell])$ by parallelizing/distributing computations across the particles (recall that each particle’s update is independent of the other particles’).
Given the definition of the metric $\mathsf{d}$ in (10), Theorem 7 bounds the $\mathcal{L}^{2}$ error of the parameter estimates $\Theta_{K}^{N,h}$ and, also, the expected $W_{2}$ distance between the particles’ empirical distribution $Q_{K}^{N,h}$ and that, $Q_{*}^{N}$, of $N$ i.i.d. samples drawn from the optimal posterior $\pi_{\theta_{*}}$. We find the latter insightful because, in many ways, $Q_{*}^{N}$ is the best one could hope to achieve using an $\mathcal{O}(N)$-cost Monte Carlo algorithm that returns $N$-sample approximations $\pi_{\theta_{*}}$. The expected $W_{2}$ distance between $Q^{N}_{*}$ itself and the optimal posterior $\pi_{\theta_{*}}$ is essentially a property of $\pi_{\theta_{*}}$ and has more to do with the fundamental limitations of approximating $\pi_{\theta_{*}}$ using $N$-sample ensembles than the method used to obtain the ensemble. To bound this distance (and, by combining this bound with Theorem 7, to obtain bounds on the distance between $Q_{K}^{N,h}$ and $\pi_{\theta_{*}}$), we refer the reader to [19, 14] and references therein.
@@
If we ‘warm start’ PGD (see, e.g., [12]) by initializing the parameter estimates and the particles at the log-likelihood $\ell$ maxima (i.e., with $\theta_{0}=0$ and $q_{0}=\delta_{0}$), $B_{0}$ in (27) vanishes and we obtain the following sharper bound.
@@
$\mathsf{d}((\Theta_{K}^{N,h},Q_{K}^{N,h}),(\theta_{*},Q_{*}^{N}))\leq\sqrt{h\frac{4h+4/\iota}{\iota}220L^{2}\left(\frac{L^{2}hd_{x}}{\lambda}+d_{x}\right)}+\sqrt{\frac{d_{x}}{\lambda}}\left(\frac{L\sqrt{2}}{\lambda\sqrt{N}}+e^{-h\lambda K}\right).$
@@
The corresponding cost is $\mathcal{O}(\epsilon^{-4}\log(M\epsilon^{-1})[\text{eval. cost of }\tilde{\ell}])$, where we are ignoring any dependence on $\tilde{d}_{x}$ which is usually small for these models. In particular, the cost depends on $M$ only through the evaluation cost of $\ell$, which cannot be avoided. We can, however, mitigate this by parellizing/distributing over both particles and datapoints, so lowering the running time to $\mathcal{O}(\epsilon^{-2}\log(\epsilon^{-1})[\text{eval. cost of }\tilde{\ell}])$.
The corollary provides some theoretical support for the idea that the PGD algorithm and its derivatives can work well for problems in which $d_{x}$ is large, even when $N$ is relatively modest, provided that the structure of the model allows the log-likelihood to be decomposed as a sum of functions of low-dimensional contributions arising from individual datapoints. This appeared to be the case in several examples explored in [23, Section 3], including a ‘generator network’ model for which $d_{x}\approx 2.5\times 10^{6}$ but good performance was obtained with only $10$ particles. The model, however, lies beyond the scope of Corollary 9 because, in that case, $f_{\theta}$ in (29) is a non-convex neural network and $\theta$ encompasses its weights. Moreover, [23]’s authors adapted PGD’s step sizes to account for this and also subsampled $\ell$’s gradient to facilitate efficient implementation on a single GPU. Regardless, we view the corollary as a tentative first step in explaining the behaviour observed in [23].
@@ 3.1 PGD’s derivation and Theorem 7’s proof
Under Assumptions 1(iii) and 35, if $q_{0}$ belongs to $\mathcal{P}_{2}(\mathbb{R}^{d_{x}})$, then (31) has a strong solution $(\theta_{t},X_{t}^{1},\dots,X_{t}^{N})_{t\geq 0}$ and the solution satisfies
@@
(32) is a standard SDE with a Lipschitz drift and diffusion coefficients, so the existence of its solutions is routine (e.g. [30, Theorem 5.2.1]). The error bound follows from a synchronous coupling argument; see Appendix B.3 for details. ∎
@@ 4 Discussion
In this paper, we theoretically validated PGD (Algorithm 1) by bounding its error for models with strongly concave log-likelihoods. In order to do this we analyzed the theoretical properties of both PGD and its continuous-time infinite-particle limit: the gradient flow (5). To study the latter, we extended certain inequalities well-known in the optimal transport and optimization literatures and several results interlinking them. In doing so, we obtain conditions weaker than strong concavity of the likelihood under which the flow converges at an exponential rate to the set $\mathcal{M}_{*}$ of pairs $(\theta_{*},\pi_{\theta_{*}})$ of marginal likelihood maximizers $\theta_{*}$ and matching posterior distributions $\pi_{\theta_{*}}$. Under the additional assumptions that the model log-likelihood is strongly concave and its gradient is Lipschitz continuous, we then proved explicit non-asymptotic error bounds for PGD.
@@
Aside from evidencing PGD’s well-foundedness, our bounds serve as a starting point for theoretical comparisons of PGD with other maximum marginal likelihood algorithms. For instance, [1] analyze the IPLA algorithm which can be viewed as a modification of PGD more amenable to standard analysis and obtain very similar scalings in $\epsilon,d_{x},d_{\theta}$ for $h,N,$ and $K$ necessary to achieve an error of $\epsilon^{-1}$ under warm-start conditions (see Section 3.6 therein) to those obtained here. The main difference between these bounds is that $N$ must be proportional to $d_{x}\epsilon^{-2}$ in our case and proportional to $d_{\theta}\epsilon^{-2}$ in the case of IPLA. Their dependence on $d_{\theta}$ stems from the extra noise term featuring IPLA’s parameter updates; see Algorithm 1 and the proof of Proposition 3 in [1]. We conjecture that the dependence on $d_{x}$ present in our results but not those of [1] simply reflects the fact that our bounds controlling the error of both the parameter estimates and the particle approximation, while theirs only control the former.
Similarly, our bounds could help settle the question of whether it is possible to ‘accelerate’ PGD and achieve faster convergence rates via the use of momentum [24] as its the case for both gradient descent [28] and the unadjusted Langevin algorithm [26]. In particular, [24, Theorem 4.1] shows that, under the xLSI, the free energy $F$ along the limiting flow of such a ‘momentum-enriched’ version of PGD converges exponentially fast to $F$’s infimum. Given Theorem 2, applying our extension of the Talagrand inequality, we can immediately prove convergence in $\mathsf{d}$-distance for the momentum-enriched PGD studied in [24]. By following arguments similar to those in Section 3.1, it might be possible to then also obtain similar results for the momentum-enriched PGD and compare the algorithms’ performances in various settings.
While our motivation for this work was understanding the theoretical properties of PGD, we find the results in Section 2 interesting in their own right and we believe they might find applications beyond the study of particle-based algorithms like PGD. For instance, for models with $\lambda$-strongly concave log-likelihoods, Theorem 4 gives us a (to the best of our knowledge) novel upper bound on the optimal marginal log-likelihood:
@@ Appendix A Proofs for Section 2
A.1 Formal derivation of de Bruijin’s identity
Using the defining equation of the gradient flow (5) and integrating by parts we compute
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\int\log(q_{t}(x))q_{t}(\operatorname{d\!}x)$ $=\int(\log(q_{t}(x))+1)\nabla_{x}\cdot\bigg{[}q_{t}(x)\nabla_{x}\log\bigg{(}\frac{q_{t}(x)}{\rho_{\theta_{t}}(x)}\bigg{)}\bigg{]}\operatorname{d\!}x$
@@ A.1 Formal derivation of de Bruijin’s identity
and
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\int\ell(\theta_{t},x)q_{t}(\operatorname{d\!}x)$ $=\int\left\langle{\nabla_{\theta}\ell(\theta_{t},x)},{\dot{\theta_{t}}}\right\rangle q_{t}(\operatorname{d\!}x)-\int\ell(\theta_{t},x)\nabla_{x}\cdot\bigg{[}q_{t}(x)\nabla_{x}\log\bigg{(}\frac{q_{t}(x)}{\rho_{\theta_{t}}(x)}\bigg{)}\bigg{]}\operatorname{d\!}x$
@@
which upon re-arranging shows that formally $(\operatorname{d\!}/\operatorname{d\!}t)F(\theta_{t},q_{t})=-I(\theta_{t},q_{t})$.
@@ A.2 Proof of Theorem 2
Here we follow some arguments from [31]. Throughout this appendix, we assume that Assumptions 13 hold. Let $(\theta,q)$ be any point in $\mathcal{M}_{2}$. As we show in Lemma 13 below,
@@
Let $(\theta,q)\in\mathcal{M}_{2}$ and let $(\theta_{t},q_{t})$ satisfy (5) and define $v_{t}(x)\mathrel{\mathop{:}}=\nabla_{x}\log(q_{t}(x)/\rho_{\theta_{t}}(x))$. [2, Theorem 8.4.7] shows that
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}W_{2}(q_{t},q)^{2}=2\int\left\langle{v_{t}(x_{1})},{x_{1}-x_{2}}\right\rangle\operatorname{d\!}\varrho(x_{1},x_{2})$ (38)
@@
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}W_{2}(q_{t},q)^{2}\leq 2\sqrt{\int\mathinner{\!\left\lVert x_{1}-x_{2}\right\rVert}^{2}\operatorname{d\!}\varrho(x_{1},x_{2})\int\mathinner{\!\left\lVert v_{t}(x)\right\rVert}^{2}q(\operatorname{d\!}x)}=2W_{2}(q_{t},q)\sqrt{\int\mathinner{\!\left\lVert v_{t}(x)\right\rVert}^{2}q(\operatorname{d\!}x)}.$
@@
Inequality (36) tells us that the sequence is Cauchy in $(\mathcal{M}_{2}^{1},\mathsf{d})$ and, consequently, in $(\mathcal{M}_{2}\mathrel{\mathop{:}}=\mathbb{R}^{d_{\theta}}\times\mathcal{P}_{2}(\mathbb{R}^{d_{x}}),\mathsf{d})$. Because both $(\mathbb{R}^{d_{\theta}},\mathsf{d}_{E})$ and $(\mathcal{P}_{2}(\mathbb{R}^{d_{x}}),\mathsf{d}_{W_{2}})$ are complete metric spaces [38, Theorem 6.18], so is $(\mathcal{M}_{2},\mathsf{d})$. Hence, $(\theta_{t_{n}},q_{t_{n}})_{n\in\mathbb{N}}$ converges to a limit $(\theta_{\infty},q_{\infty})$ in $\mathcal{M}_{2}$. As we show in Lemma 15 below, the free energy $F$ is lower semicontinuous on $(\mathcal{M}_{2},\mathsf{d})$. Consequently, Theorem 1 implies that
@@
$\mathrm{KL}(q||\pi)=\sup\left\{\int\phi(x)q(\operatorname{d\!}x)-\int(e^{\phi(x)}-1)\pi(\operatorname{d\!}x)\mathrel{\mathop{:}}\phi\in\mathcal{C}_{b}(\mathbb{R}^{d_{x}})\right\},$
where $\mathcal{C}_{b}(\mathbb{R}^{d_{x}})$ denotes the set of continuous bounded functions $\mathbb{R}^{d_{x}}$. Consequently,
$F(\theta,q)=\mathrm{KL}(q||\pi_{\theta})-\log(Z_{\theta})=\sup\{G(\theta,q,\phi)\mathrel{\mathop{:}}\phi\in\mathcal{C}_{b}(\mathbb{R}^{d_{x}})\}-\log(Z_{\theta}),$
where $G(\theta,q,\phi)\mathrel{\mathop{:}}=\int\phi(x)q(\operatorname{d\!}x)-\int(e^{\phi(x)}-1)\pi_{\theta}(\operatorname{d\!}x)$. Because Assumption 1(i) implies that $\theta\mapsto\log(Z_{\theta})$ is continuous and the pointwise supremum of a family of lower semicontinuous (l.s.c.) functions is l.s.c., we need only show that $(\theta,q)\mapsto G(\theta,q,\phi)$ is l.s.c. for each $\phi$ in $\mathcal{C}_{b}(\mathbb{R}^{d_{x}})$. Since the Wasserstein-2 topology is finer than the weak topology, it suffices to show that $(\theta,q)\mapsto G(\theta,q,\phi)$ is continuous in the weak topology for any given $\phi$ in $\mathcal{C}_{b}(\mathbb{R}^{d_{x}})$. Consider a sequence $(\theta_{n},q_{n})_{n\in\mathbb{N}}$ such that $\theta_{n}\rightarrow\theta$ and $q_{n}\rightarrow q$ weakly. By definition of weak convergence, $\int\phi(x)q_{n}(\operatorname{d\!}x)\rightarrow\int\phi(x)q(\operatorname{d\!}x)$. Furthermore, since $x\mapsto(e^{\phi(x)}-1)\in\mathcal{C}_{b}(\mathbb{R}^{d_{x}})$, we have $|\int(e^{\phi(x)}-1)\pi_{\theta_{n}}(\operatorname{d\!}x)-\int(e^{\phi(x)}-1)\pi_{\theta}(\operatorname{d\!}x)|\leq c_{\phi}\int|\pi_{\theta_{n}}(x)-\pi_{\theta}(x)|\operatorname{d\!}x$ for some $c_{\phi}<\infty$. Since $\theta\mapsto\pi_{\theta}$ is continuous by Assumption 1(i), an application of Scheffé’s lemma concludes the proof. ∎

### 11 Apr 2024 v2

@@ 2.2 The Talagrand inequality and flow’s exponential convergence
In this case, Theorem 2 reduces to the well-known result stating that the QG condition holds whenever the PŁI (21) holds; see [23, Theorem 2].
@@ 2.3 Strongly log-concave models
Any model satisfying Assumptions 1(iii) and 3 satisfies the xLSI.
@@
Given that (11) with $q=\pi_{\theta}$ reduces to the PŁI (21) with $f(\theta)=-\log(Z_{\theta})$, Theorem 4 extends the well-known fact that the PŁI holds for strongly convex $f$; e.g., see [23, Theorem 2].
@@
Under Assumption 3, the marginal likelihood has a unique maximizer $\theta_{*}$; e.g., see [26, Theorem 4 in Appendix B.3]. Hence, Corollary 3 and (24) tell us that the flow converges exponentially fast to $(\theta_{*},\pi_{\theta_{*}})$:
@@
If Assumptions 1(iii), 23 hold and $(\theta_{0},q_{0})$ belongs to $\mathcal{M}_{2}$, then
@@ 3 Error bounds for particle gradient descent
If Assumptions 1(ii) and 4 hold and $(\theta,X)$ belongs to $\mathbb{R}^{d_{\theta}}\times\mathcal{L}^{2}(\mathbb{R}^{d_{x}})$, then the SDE (6) has an unique strong solution $(\theta_{t},X_{t})_{t\geq 0}$ such that $(\theta_{0},X_{0})=(\theta,X)$. Moreover, $(\theta_{t},\textup{Law}(X_{t}))_{t\geq 0}$ is a classical solution of the gradient flow (5) satisfying Assumption 2.
@@
Under Assumptions 1(ii–iii) and 34, if $X_{0}^{1,h},\dots,X_{0}^{N,h}$ are drawn independently from a distribution $q_{0}$ in $\mathcal{P}_{2}(\mathbb{R}^{d_{x}})$ and $h\leq 1/(\lambda+L)$, then
$\mathsf{d}((\Theta_{K}^{N,h},Q_{K}^{N,h}),(\theta_{*},Q_{*}^{N}))\leq\sqrt{h}A_{0,h}+\frac{L\sqrt{2}}{\lambda\sqrt{N}}\sqrt{B_{0}+\frac{2d_{x}}{\lambda}}+\mathsf{d}((\theta_{0},q_{0}),(\theta_{*},\pi_{\theta_{*}}))e^{-h\lambda K}$ (27)
@@
We postpone the proof until Section 3.1 below. In short, the Theorem 7 gives us the bound (9) and an explicit expression for the proportionality constant as a function of the algorithm’s parameters, the model’s Lipschitz and concavity constants, the initial condition, and the dimension of the latent space. It follows that in order to achieve an error of $\mathcal{O}(\epsilon)$, we need only choose the step size $h$ proportional to $\epsilon^{2}$, the particle number $N$ proportional to $\epsilon^{-2}$, and the step number $K$ proportional to $\epsilon^{-2}\log(\epsilon^{-1})$. The corresponding computational cost is $\mathcal{O}(\epsilon^{-4}\log(\epsilon^{-1})[\text{eval. cost of }\ell])$; and we can lower the running time to $\mathcal{O}(\epsilon^{-2}\log(\epsilon^{-1})[\text{eval. cost of }\ell])$ by parallelizing/distributing computations across the particles (recall that each particle’s update is independent of the other particles’). Given the definition of the metric $\mathsf{d}$ in (10), Theorem 7 bounds the $\mathcal{L}^{2}$ error of the parameter estimates $\Theta_{K}^{N,h}$ and, also, the expected $W_{2}$ distance between the particles’ empirical distribution $Q_{K}^{N,h}$ and that, $Q_{*}^{N}$, of $N$ i.i.d. samples drawn from the optimal posterior $\pi_{\theta_{*}}$. We find the latter insightful because, in many ways, $Q_{*}^{N}$ is the best one could hope to achieve using an $\mathcal{O}(N)$-cost Monte Carlo algorithm that returns $N$-sample approximations $\pi_{\theta_{*}}$. The expected $W_{2}$ distance between $Q^{N}_{*}$ itself and the optimal posterior $\pi_{\theta_{*}}$ is essentially a property of $\pi_{\theta_{*}}$ and has more to do with the fundamental limitations of approximating $\pi_{\theta_{*}}$ using $N$-sample ensembles than the method used to obtain the ensemble. To bound this distance (and, by combining this bound with Theorem 7, to obtain bounds on the distance between $Q_{K}^{N,h}$ and $\pi_{\theta_{*}}$), we refer the reader to [21, 16] and references therein.
@@
If we ‘warm start’ PGD (see, e.g., [14]) by initializing the parameter estimates and the particles at the log-likelihood $\ell$ maxima (i.e., with $\theta_{0}=0$ and $q_{0}=\delta_{0}$), then $B_{0}$ in (27) vanishes and we obtain the following sharper bound.
@@
$\mathsf{d}((\Theta_{K}^{N,h},Q_{K}^{N,h}),(\theta_{*},Q_{*}^{N}))\leq\sqrt{h\frac{4h+4/\iota}{\iota}220L^{2}\left(\frac{2L^{2}hd_{x}}{\lambda}+d_{x}\right)}+\sqrt{\frac{d_{x}}{\lambda}}\left(\frac{2L}{\lambda\sqrt{N}}+e^{-h\lambda K}\right).$
@@
The corresponding cost is $\mathcal{O}(\epsilon^{-4}\log(\epsilon^{-1})[\text{eval. cost of }\tilde{\ell}])$, where we are ignoring any dependence on $\tilde{d}_{x}$ which is usually small for these models. In particular, the cost depends on $M$ only through the evaluation cost of $\ell$, which cannot be avoided. We can, however, mitigate this by parellizing/distributing over both particles and datapoints, so lowering the running time to $\mathcal{O}(\epsilon^{-2}\log(\epsilon^{-1})[\text{eval. cost of }\tilde{\ell}])$. The corollary provides some theoretical support for the idea that the PGD algorithm and its derivatives can work well for problems in which $d_{x}$ is large, even when $N$ is relatively modest, provided that the structure of the model allows the log-likelihood to be decomposed as a sum of functions of low-dimensional contributions arising from individual datapoints. This appeared to be the case in several examples explored in [26, Section 3], including a ‘generator network’ model for which $d_{x}\approx 2.5\times 10^{6}$ but good performance was obtained with only $10$ particles. The model, however, lies beyond the scope of Corollary 9 because, in that case, $f_{\theta}$ in (29) is a non-convex neural network and $\theta$ encompasses its weights. Moreover, [26]’s authors adapted PGD’s step sizes to account for this and also subsampled $\ell$’s gradient to facilitate efficient implementation on a single GPU. Regardless, we view the corollary as a tentative first step in explaining the behaviour observed in [26].
@@ 3.1 PGD’s derivation and Theorem 7’s proof
Under Assumptions 1(ii–iii) and 34, if $q_{0}$ belongs to $\mathcal{P}_{2}(\mathbb{R}^{d_{x}})$, then (31) has a strong solution $(\theta_{t},X_{t}^{1},\dots,X_{t}^{N})_{t\geq 0}$ and the solution satisfies
@@
(32) is a standard SDE with a Lipschitz drift and diffusion coefficients, so the existence of its solutions is routine (e.g. [35, Theorem 5.2.1]). The error bound follows from a synchronous coupling argument; see Appendix B.3 for details. ∎
@@ 4 Discussion
In this paper, we theoretically validated PGD (Algorithm 1) by bounding its error for models with strongly concave log-likelihoods. In order to do this we analyzed the theoretical properties of both PGD and its continuous-time infinite-particle limit: the gradient flow (5). To study the latter, we extended certain inequalities well-known in the optimal transport and optimization literatures and several results interlinking them. In doing so, we obtain conditions weaker than strong concavity of the likelihood under which the flow converges at an exponential rate to the set $\mathcal{M}_{*}$ of pairs $(\theta_{*},\pi_{\theta_{*}})$ of marginal likelihood maximizers $\theta_{*}$ and matching posterior distributions $\pi_{\theta_{*}}$. Under the additional assumptions that the model log-likelihood is strongly concave and its gradient is Lipschitz continuous, we then proved explicit non-asymptotic error bounds for PGD. Aside from evidencing PGD’s well-foundedness, our bounds serve as a starting point for theoretical comparisons of PGD with other maximum marginal likelihood algorithms. For instance, [2] analyze the IPLA algorithm which can be viewed as a modification of PGD more amenable to standard analysis and obtain very similar scalings in $\epsilon,d_{x},d_{\theta}$ for $h,N,$ and $K$ necessary to achieve an error of $\epsilon^{-1}$ under warm-start conditions (see Section 3.6 therein) to those obtained here. The main difference between these bounds is that $N$ must be proportional to $d_{x}\epsilon^{-2}$ in our case and proportional to $d_{\theta}\epsilon^{-2}$ in the case of IPLA. Their dependence on $d_{\theta}$ stems from the extra noise term featuring IPLA’s parameter updates; see Algorithm 1 and the proof of Proposition 3 in [2]. We conjecture that the dependence on $d_{x}$ present in our results but not those of [2] simply reflects the fact that our bounds controlling the error of both the parameter estimates and the particle approximation, while theirs only control the former. Similarly, our bounds could help settle the question of whether it is possible to ‘accelerate’ PGD and achieve faster convergence rates via the use of momentum [28] as its the case for both gradient descent [33] and the unadjusted Langevin algorithm [30]. In particular, [28, Theorem 4.1] shows that, under the xLSI, the free energy $F$ along the limiting flow of such a ‘momentum-enriched’ version of PGD converges exponentially fast to $F$’s infimum. Given Theorem 2, applying our extension of the Talagrand inequality, we can immediately prove convergence in $\mathsf{d}$-distance for the momentum-enriched PGD studied in [28]. By following arguments similar to those in Section 3.1, it might be possible to then also obtain similar results for the momentum-enriched PGD and compare the algorithms’ performances in various settings. While our motivation for this work was understanding the theoretical properties of PGD, we find the results in Section 2 interesting in their own right and we believe they might find applications beyond the study of particle-based algorithms like PGD. For instance, for models with $\lambda$-strongly concave log-likelihoods, Theorem 4 gives us a (to the best of our knowledge) novel upper bound on the optimal marginal log-likelihood:
@@
Acknowledgements
JK and AMJ acknowledge support from the Engineering and Physical Sciences Research Council (EPSRC; grant EP/T004134/1). AMJ acknowledges further support from the EPSRC (grants EP/R034710/1 and EP/Y014650/1). RC was funded by the UK Engineering and Physical Sciences Research Council (EPSRC) via studentship 2585619 as part of grant number EP/W523793/1.
@@ Appendix A Proofs for Section 2
A.1 de Bruijin’s identity
Lemma 13.
If Assumption 2 holds, then also equation (17) does.
Proof.
The regularity of $(\theta_{t},q_{t})_{t\geq 0}$ allow us to exchange differentiation and integration by the Leibniz rule and compute
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\int\log(q_{t}(x))q_{t}(\operatorname{d\!}x)$ $=\int(\log(q_{t}(x))+1)\nabla_{x}\bigg{[}q_{t}(x)\nabla_{x}\log\bigg{(}\frac{q_{t}(x)}{\rho_{\theta_{t}}(x)}\bigg{)}\bigg{]}\operatorname{d\!}x$
@@
where we integrated by parts, and
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\int\ell(\theta_{t},x)q_{t}(\operatorname{d\!}x)$ $=\int\left\langle{\nabla_{\theta}\ell(\theta_{t},x)},{\dot{\theta_{t}}}\right\rangle q_{t}(\operatorname{d\!}x)-\int\ell(\theta_{t},x)\nabla_{x}\bigg{[}q_{t}(x)\nabla_{x}\log\bigg{(}\frac{q_{t}(x)}{\rho_{\theta_{t}}(x)}\bigg{)}\bigg{]}\operatorname{d\!}x$
@@
Upon re-arranging, this shows $(\operatorname{d\!}/\operatorname{d\!}t)F(\theta_{t},q_{t})=-I(\theta_{t},q_{t})$. ∎
@@ A.2 Proof of Theorem 2
Here some arguments are similar to those in [36]. Throughout this appendix, we assume that Assumptions 12 hold. Let $(\theta,q)$ be any point in $\mathcal{M}_{2}$. As we show in Lemma 14 below,
@@
Let $(\theta,q)\in\mathcal{M}_{2}$, take $(\theta_{t},q_{t})_{t\geq 0}$ as per Assumption 2 and define the velocity field $v_{t}(x)\mathrel{\mathop{:}}=\nabla_{x}\log(q_{t}(x)/\rho_{\theta_{t}}(x))$. [3, Theorem 8.4.7] shows that
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\mathsf{d}_{W_{2}}(q_{t},q)^{2}=2\int\left\langle{v_{t}(x_{1})},{x_{1}-x_{2}}\right\rangle\operatorname{d\!}\varrho(x_{1},x_{2})$ (38)
@@
$\frac{\operatorname{d\!}}{\operatorname{d\!}t}\mathsf{d}_{W_{2}}(q_{t},q)^{2}\leq 2\sqrt{\int\mathinner{\!\left\lVert x_{1}-x_{2}\right\rVert}^{2}\operatorname{d\!}\varrho(x_{1},x_{2})\int\mathinner{\!\left\lVert v_{t}(x)\right\rVert}^{2}q(\operatorname{d\!}x)}=2\mathsf{d}_{W_{2}}(q_{t},q)\sqrt{\int\mathinner{\!\left\lVert v_{t}(x)\right\rVert}^{2}q(\operatorname{d\!}x)}.$
@@
Inequality (36) tells us that the sequence is Cauchy in $(\mathcal{M}_{2}^{1},\mathsf{d})$ and, consequently, in $(\mathcal{M}_{2}\mathrel{\mathop{:}}=\mathbb{R}^{d_{\theta}}\times\mathcal{P}_{2}(\mathbb{R}^{d_{x}}),\mathsf{d})$. Because both $(\mathbb{R}^{d_{\theta}},\mathsf{d}_{E})$ and $(\mathcal{P}_{2}(\mathbb{R}^{d_{x}}),\mathsf{d}_{W_{2}})$ are complete metric spaces [45, Theorem 6.18], so is $(\mathcal{M}_{2},\mathsf{d})$. Hence, $(\theta_{t_{n}},q_{t_{n}})_{n\in\mathbb{N}}$ converges to a limit $(\theta_{\infty},q_{\infty})$ in $\mathcal{M}_{2}$. As we show in Lemma 16 below, the free energy $F$ is lower semicontinuous on $(\mathcal{M}_{2},\mathsf{d})$. Consequently, Theorem 1 implies that
@@
$\mathrm{KL}(q||\pi)=\sup\left\{\int\phi(x)q(\operatorname{d\!}x)-\int(e^{\phi(x)}-1)\pi(\operatorname{d\!}x)\mathrel{\mathop{:}}\phi\in\mathcal{C}_{b}(\mathbb{R}^{d_{x}},\mathbb{R})\right\},$
where $\mathcal{C}_{b}(\mathbb{R}^{d_{x}},\mathbb{R})$ denotes the set of real valued continuous bounded functions $\mathbb{R}^{d_{x}}$. Consequently,
$F(\theta,q)=\mathrm{KL}(q||\pi_{\theta})-\log(Z_{\theta})=\sup\{G(\theta,q,\phi)\mathrel{\mathop{:}}\phi\in\mathcal{C}_{b}(\mathbb{R}^{d_{x}},\mathbb{R})\}-\log(Z_{\theta}),$
where $G(\theta,q,\phi)\mathrel{\mathop{:}}=\int\phi(x)q(\operatorname{d\!}x)-\int(e^{\phi(x)}-1)\pi_{\theta}(\operatorname{d\!}x)$. Because Assumption 1(i) implies that $\theta\mapsto\log(Z_{\theta})$ is continuous and the pointwise supremum of a family of lower semicontinuous (l.s.c.) functions is l.s.c., we need only show that $(\theta,q)\mapsto G(\theta,q,\phi)$ is l.s.c. for each $\phi$ in $\mathcal{C}_{b}(\mathbb{R}^{d_{x}},\mathbb{R})$. Since the Wasserstein-2 topology is finer than the weak topology, it suffices to show that $(\theta,q)\mapsto G(\theta,q,\phi)$ is continuous in the weak topology for any given $\phi$ in $\mathcal{C}_{b}(\mathbb{R}^{d_{x}},\mathbb{R})$. Consider a sequence $(\theta_{n},q_{n})_{n\in\mathbb{N}}$ such that $\theta_{n}\rightarrow\theta$ and $q_{n}\rightarrow q$ weakly. By definition of weak convergence, $\int\phi(x)q_{n}(\operatorname{d\!}x)\rightarrow\int\phi(x)q(\operatorname{d\!}x)$. Furthermore, since $x\mapsto(e^{\phi(x)}-1)\in\mathcal{C}_{b}(\mathbb{R}^{d_{x}},\mathbb{R})$, we have $|\int(e^{\phi(x)}-1)\pi_{\theta_{n}}(\operatorname{d\!}x)-\int(e^{\phi(x)}-1)\pi_{\theta}(\operatorname{d\!}x)|\leq c_{\phi}\int|\pi_{\theta_{n}}(x)-\pi_{\theta}(x)|\operatorname{d\!}x$ for some $c_{\phi}<\infty$. Since $\theta\mapsto\pi_{\theta}$ is continuous by Assumption 1(i), an application of Scheffé’s lemma concludes the proof. ∎